find the sum of the following sequence: 3, 7, 11, 15,..., 79?
Found 2 solutions by josgarithmetic, stanbon:
Answer by josgarithmetic(39618)
Answer by stanbon(75887)

find the sum of the following sequence: 3, 7, 11, 15,..., 79?
Determine "n"::
79 = 3 + (n-1)4
76 = (n-1)4
n-1 = 19
n = 20
----------------
Sum = (n/2)(a(1) + a(20))
Sum = 10(3+79) = 10*82 = 820
-----------------
Cheers,
Stan H.
--------------